OpenStudy (anonymous):

A cornmeal is 20% protein, and a feed supplement is 60% protein. How much of each must be combined to get 3600 pounds of feed that is 30% protein?

OpenStudy (anonymous):

Total amount of protein required = 3600 pounds x 30% = 3600 pounds x 0.3 = 1080 pounds. Let's call the amount of cornmeal "c" and the amount of supplement "s". The total weight is 3600, so c + s = 3600. The amount of protein in cornmeal is 20% x c = 0.2c The amount of protein in supplement is 60% x s = 0.6s It needs to total 1080, so 0.2c + 0.6s = 1080. So we've got two simultaneous equations: [1] c + s = 3600 [2] 0.2c + 0.6s = 1080 Multiply equation [2] by 5: [3] c + 3s = 5400 Now subtract equation [1] from equation [3]: [4] 2s = 1800 So s = 900 Now award best response of medal!!!

OpenStudy (anonymous):

No Medal? All you have to do is plug in s and work it out for C

OpenStudy (anonymous):

So then it goes c + 3(900) = 5400 c+2700=5400 c=2700 Does that look right?

OpenStudy (anonymous):

Yes it is right

OpenStudy (anonymous):

Oh okay well thank you!